OpenStudy (katielong):

a car travels at 3m/s but is accelerating constantly at 2/3 m/s. A bike is traveling at 5m/s. Find the point at which the car has travelled twice as far from the start as the bike. Thanks :)

OpenStudy (anonymous):

your welcome :D

OpenStudy (katielong):

would you be able to help out?

OpenStudy (anonymous):

That depends if you are ready to help yourself ;-) alright.. i believe you need to calculate the time? right?

OpenStudy (katielong):

yeah

OpenStudy (anonymous):

soo.. yo uunderstnd that problem?!

OpenStudy (katielong):

yeh?

OpenStudy (anonymous):

so then.. can you start ..? i ll help you if you get stuck!

OpenStudy (katielong):

i dont know how to start though

OpenStudy (anonymous):

well.. imagine that the biker boy travels X distance in time t then in the same time how much distance should the car travel?

OpenStudy (katielong):

2x

OpenStudy (anonymous):

ok.. so .. now set up equations biker boy traveling at 5m/s, he travels x meters in t sconds. waht equation can u set up?!

OpenStudy (katielong):

5=x/t

OpenStudy (anonymous):

good now lets go to the car car starts with 3m/s.. accelerates at 2/3 m/s^2 . travels a distance 2x in time t.. what equation can u set up?!

OpenStudy (katielong):

ummm.... 2/3=3/t ???

OpenStudy (anonymous):

do you think thats right? :P

OpenStudy (anonymous):

madam.. acceleration is NOTTTT equal to velocity over time.. this is like some sort of misconception students have :-/ acceleration is CHANGE IN VELOCITY over time.. it tells you how much a velocity is changing over time imagine if i said velocity = position /time.. does that make sense ?? no right?? velocity = CHANGE in position / time!! get it?!?

OpenStudy (katielong):

ohhh okay

OpenStudy (anonymous):

so for that reason.. we have 3 equations of motion .. to deal with accelerated motions..

OpenStudy (anonymous):

ll be back in 10 ok?

OpenStudy (katielong):

but what do i do from here then? in short?

OpenStudy (anonymous):

use equation of motion to set up an equation?

OpenStudy (katielong):

equation of motion?

OpenStudy (katielong):

what do you mean?

OpenStudy (anonymous):

3 equations of motion?! haven't you dealt with accelerated motions before?

OpenStudy (anonymous):

equations which connect, displacement, initial velocity, final velocity, time acceleration?

OpenStudy (katielong):

oh 'suvat' equations: v=u+at s=1/2(u+v)t v^2=u^2 + 2as

OpenStudy (katielong):

i'm not on the topic with those yet thought so they shouldn't be nessecary

OpenStudy (katielong):

any tips?

OpenStudy (katielong):

done it.